Pamela owns 36% of the stock of a corporation and sells 2/3 of her stock at a 10% profit. What percent of this stock does she own after this transaction?
Answer by adamchapman(301)

36 - (36 x (2/3)) = 36 - 24 = 12
Pamela now owns 12% of the companys stock
I hope this helps.
Adam
P.S. please visit my website, I only set it up last week and it may be of use to you. The address is http://geocities.com/quibowibbler